# Symplectic mapping class relations from pencil pairs

### v1

@@ 5. $\dim_{\mathbb{C}}X^{\pm}=3$ examples
Now we work out some new examples of positive mapping class relations coming from $\dim_{\mathbb{C}}X^{\pm}=3$ pencils by applying the results of §4. First, we briefly review some vocabulary and known results.
@@
    its index, $\operatorname{ind}_{X}$ is the maximal $k$ for which there is an ample $\operatorname{\mathcal{L}}\rightarrow X$ such that $A_{X}=\operatorname{\mathcal{L}}^{\otimes k}$, and
@@
A simply connected $\dim_{\mathbb{C}}=2$ Calabi-Yau variety is a $K_{3}$ surface. According to Kodaira, the diffeomorphism type is uniquely determined. See [H16] for this and additional background. For an ample line bundle $\operatorname{\mathcal{L}}$ over a $K_{3}$ surface $Z$, the genus is define as
$g(Z,\operatorname{\mathcal{L}})=1+\frac{1}{2}\int_{Z}c_{1}(\operatorname{\mathcal{L}})^{2}$
and computes the genus of a hypersurface in $Z$ cut out by a section of $\operatorname{\mathcal{L}}$. It follows a Hodge theory computation that $\chi(Z)=24$. The adjunction formula applied to anticanonical divisors that for a $\dim_{\mathbb{C}}X=3$ Fano $\int_{X}c_{1}(X)c_{2}(X)=24$ as well, so that the only interesting Chern numbers of a Fano $X$ are $\int_{X}c_{1}(X)^{3}=\int_{X}A_{X}^{3}$ and $\chi(X)$. Combining these observations with some arithmetic of Chern classes, we obtain the following lemma.
@@
If $(\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1})$ is a pencil associated to a $\operatorname{\mathcal{L}}$ on a $\dim_{\mathbb{C}}X=3$ variety has
#$\#\operatorname{Crit}\pi=-\chi(X)+\int_{X}2c_{2}(X)c_{1}(\operatorname{\mathcal{L}})-3c_{1}(X)c_{1}(\operatorname{\mathcal{L}})^{2}+2c_{1}(\operatorname{\mathcal{L}})^{3}.$
So when $\operatorname{\mathcal{L}}=A_{X}^{\otimes k}$ on a Fano $X$,
#$\#\operatorname{Crit}\pi=-\chi(X)+48k+k^{2}(2k-3)\int_{X}A_{X}^{3}.$
5.1. Classification of polarized $K_{3}$ surfaces of low genus
We will be interested in $\dim_{\mathbb{C}}X=3$ Fano varieties for which we have a very ample $\operatorname{\mathcal{L}}\rightarrow X$ satisfying $\operatorname{\mathcal{L}}^{\otimes\operatorname{ind}_{X}}=A_{X}$. This implies that $\operatorname{\mathcal{L}}|_{Z}$ is very ample as well for $Z\subset X$ an anticanonical divisor. The following theorem reviews classification results of such polarized $K_{3}$ surfaces $(Z,\operatorname{\mathcal{L}})$ by Saint-Donat [SD74] (for $g(Z,\operatorname{\mathcal{L}})\leq 5$) and Mukai [M87] (for $6\leq g(Z,\operatorname{\mathcal{L}})\leq 10$).
@@ 5.1. Classification of polarized $K_{3}$ surfaces of low genus
$g=1+\frac{1}{2}\int_{Z}c_{1}(\operatorname{\mathcal{L}})^{2}=1+\frac{1}{2}\operatorname{ind}_{X}\int_{X}c_{1}(\operatorname{\mathcal{L}})^{3}=1+\frac{1}{2}\operatorname{ind}_{X}^{-2}\int_{X}A_{X}^{3}.$
Then $(Z,\operatorname{\mathcal{L}})$ has a projective model $(\mathcal{M}_{g}\subset\mathbb{P}^{N_{g}},\deg_{g})$ which is entirely determined by $g$.
Note that we require $\operatorname{\mathcal{L}}$ to be very ample, rather than just ample which is insufficient for the desired result [M87]. For $g=3$, $Z$ is a $\deg=4$ hypersurface in $\mathbb{P}^{3}$ with $\operatorname{\mathcal{L}}=\mathcal{O}_{1}$ so the model is $(\mathcal{M}_{3}=\mathbb{P}^{4}\subset\mathbb{P}^{4},\deg_{4}=(1,4))$. For $g=4$, $Z$ is an intersection of $\deg=2$ and $\deg=3$ hypersurfaces in $\mathbb{P}^{4}$ with $\operatorname{\mathcal{L}}=\mathcal{O}_{1}\rightarrow\mathbb{P}^{4}$, so the model is $(\mathcal{M}_{5}=\mathbb{P}^{5},\deg_{5}=(1,2,3))$. For $g=5$, $Z$ is an intersection of three $\deg=2$ hypersurfaces in $\mathbb{P}^{5}$ with $\operatorname{\mathcal{L}}=\mathcal{O}_{1}\rightarrow\mathbb{P}^{5}$, so $(\mathcal{M}_{5}=\mathbb{P}^{5},\deg_{5}=(1,2,2,2))$. In the $6\leq g\leq 10$ cases, the $\mathcal{M}_{g}$ have positive codimension in $\mathbb{P}^{N_{g}}$ and will be described in the following subsection.
@@ 5.2. Examples of pencil pairs from anticanonical divisors on Fano $3$-folds
Combining the above results with Lemmas 4.1.2 and 4.2.1, we have the following theorem.
@@
To find all examples of pencil pairs determinable from Theorem 5.2.1, we cross reference the above theorems with the complete classification of $\dim_{\mathbb{C}}=3$ Fano varieties by Iskovskih [I77, I77] and Mori-Mukai [MM82]. There are $105$ such varieties which are listed in the Fanography database [Be24] along with relevant data such as indices and Hodge diamonds (which of course determine Euler characteristics). [^1] See https://www.fanography.info/about for a complete list of resources from which the database is compiled.
We collect the relevant examples of our Fano $X$ into families (having at least two members) organized by their $\operatorname{ind}_{X}$ and $\int_{X}A_{X}^{3}$. Members of the families are labeled by their indices $a$-$b$ as in [Be24], with $a$ being the Picard rank. For each, a short description of $X$ is given. According to Theorem 5.2.1 any two members of a given family form a pencil pair.
5.2.1. Case $\operatorname{ind}_{X}=1,\int_{X}A_{X}^{3}=10\implies g(Z,A_{X})=6$
@@
    $X$ is an intersection of the $\mathcal{M}_{6}=G(2,5)\subset\mathbb{P}^{9}$ by two hyperplanes and a quadric having $\chi=-16$.
@@
    $X$ is a divisor of $\deg=(1,3)$ on $\mathbb{P}^{1}\times\mathbb{P}^{3}$ having $\chi=-14$.
Here is the projective model is $(\mathcal{M}_{6},\deg=(1,1,1,2))$ following [M87].
We obtain Weinstein $4$-manifolds $(F_{k},\beta_{k})$ given by a $A_{X}^{\otimes k}$ divisor in either $X$, minus its intersection with another such generic divisor. Applying Theorem 5.2.1, we have factorizations of $\tau_{\partial F_{k}}$ into products of
$\chi+48k+10k^{2}(2k-3),\quad\chi=-16,-14$
@@
5.2.2. Case $\operatorname{ind}_{X}=1,\int_{X}A_{X}^{3}=12\implies g(Z,A_{X})=7$
  1. (1-6)
    $X$ is an intersection of a $\mathcal{M}_{7}\subset\mathbb{P}^{15}$ of $\dim_{\mathbb{C}}\mathcal{M}=10$ with four hyperplanes having $\chi(X)=-10$ as described in [M87].
  2. (2-5)
    $X$ is an intersection of $\deg=(0,3)$ and $\deg=(1,1)$ hypersurfaces in $\mathbb{P}^{1}\times\mathbb{P}^{4}$ having $\chi(X)=-6$.
  3. (2-6)
    $X$ is a $\deg=(2,2)$ divisor in $\mathbb{P}^{2}\times\mathbb{P}^{2}$ having $\chi(X)=-12$.
  4. (3-1)
    $X$ is a double branched cover of $(\mathbb{P}^{1})^{3}$ over a $\deg=(2,2,2)$ divisor having $\chi(X)=-8$.
In this case the projective model of [M87] is $(\mathcal{M}_{7},(1,1,1,1,1))$.
$-\chi+48k+12k^{2}(2k-3),\quad\chi=-10,-6,-12,-8$
Dehn twists along exact Lagrangian spheres.
5.2.3. Case $\operatorname{ind}_{X}=1,\int_{X}A_{X}^{3}=14\implies g(Z,A_{X})=8$
  1. (1-7)
    $X$ is a the intersection of $\mathcal{M}_{8}=G(2,6)\subset\mathbb{P}^{14}$ with nine hyperplanes having $\chi(X)=-6$.
  2. (2-7)
    $X$ is an intersection of $\deg=(0,2)$ and $\deg=(1,2)$ divisors on $\mathbb{P}^{1}\times\mathbb{P}^{4}$ having $\chi(X)=-4$.
  3. (2-8)
    $X$ is a double branched cover of $\mathbb{P}(\mathbb{C}\oplus\mathcal{O}_{1}\rightarrow\mathbb{P}^{2})$ over an anticanonical divisor with $\chi(X)=-12$.
  4. (3-2)
    $X$ is a divisor in a $\mathbb{P}^{2}$ bundle over $\mathbb{P}^{2}$ having $\chi(X)=2$.
$-\chi+48k+14k^{2}(2k-3),\quad\chi=-6,-4,-12,2$
Dehn twists along exact Lagrangian spheres.
5.2.4. Case $\operatorname{ind}_{X}=1,\int_{X}A_{X}^{3}=16\implies g(Z,A_{X})=9$

### v2

@@ 5. $\dim_{\mathbb{C}}X^{\pm}=3$ examples
Now we work out some new examples of positive mapping class relations coming from $\dim_{\mathbb{C}}X^{\pm}=3$ pencil pairs by applying the results of §4. First, we briefly review some vocabulary and known results.
@@
    $\operatorname{ind}_{X}$ is the maximal $k$ for which there is an ample line bundle $\operatorname{\mathcal{L}}\to X$ with $A_{X}=\operatorname{\mathcal{L}}^{\otimes k}$, and
@@
A simply connected $\dim_{\mathbb{C}}=2$ Calabi-Yau variety is a $K3$ surface. According to Kodaira, the diffeomorphism type of a $K3$ surface is uniquely determined. See [H16] for this and additional background. For an ample line bundle $\operatorname{\mathcal{L}}$ over a $K3$ surface $Z$, the genus is defined as
$g(Z,\operatorname{\mathcal{L}})=1+\frac{1}{2}\int_{Z}c_{1}(\operatorname{\mathcal{L}})^{2},$
which computes the genus of a divisor in $Z$ cut out by a generic holomorphic section of $\operatorname{\mathcal{L}}$. It follows from a Hodge theory computation that $\chi(Z)=24$. The adjunction formula applied to anticanonical divisors for a $\dim_{\mathbb{C}}X=3$ Fano variety then gives
(5.0.1) $\int_{X}c_{1}(X)c_{2}(X)=24$
as well. Therefore the only interesting Chern numbers of a Fano $3$-fold $X$ are
$\int_{X}c_{1}(X)^{3}=\int_{X}c_{1}(A_{X})^{3},\quad\int_{X}c_{2}(X)=\chi(X).$
Combining these observations with some arithmetic of Chern classes, we obtain the following lemma.
@@
If $(\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1})$ is a pencil associated to a line bundle $\operatorname{\mathcal{L}}$ over a variety $X$ of $\dim_{\mathbb{C}}X=3$, then the associated Lefschetz fibration on $X\setminus\{\boldsymbol{\mathfrak{s}}_{0}=0\}$ has
#$\#\operatorname{Crit}\pi=-\chi(X)+\int_{X}2c_{2}(X)c_{1}(\operatorname{\mathcal{L}})-3c_{1}(X)c_{1}(\operatorname{\mathcal{L}})^{2}+4c_{1}(\operatorname{\mathcal{L}})^{3}$
critical points. In the specific case that $\operatorname{\mathcal{L}}=A_{X}^{\otimes k}$ over a Fano variety $X$,
#$\#\operatorname{Crit}\pi=-\chi(X)+48k+k^{2}(4k-3)\int_{X}c_{1}(A_{X})^{3}.$
Proof.
To count #$\#\operatorname{Crit}\pi$ we seek to apply the formula from Theorem 2.4.2, which requires computations of $\chi(Z_{1})$ and $\chi(B)$ where $Z_{1}$ is a divisor of $X$ cut out by some generic section $\boldsymbol{\mathfrak{s}}_{1}$ of $\operatorname{\mathcal{L}}\to X$ and $B$ is the base locus of a pencil. Throughout we apply axiomatic properties of Chern classes, cf. [MiS74].
Since the $\operatorname{\mathcal{L}}$-valued $1$-form $d\boldsymbol{\mathfrak{s}}_{1}$ is non-degenerate along $Z_{1}$, the coimage of the map $TX|_{Z_{1}}\to\operatorname{\mathcal{L}}|_{Z_{1}}$, $v_{z}\mapsto d\boldsymbol{\mathfrak{s}}_{1}(v_{z})$ is a normal bundle to $Z_{1}\subset X$. Thus we have an identification $TX|_{Z_{1}}\simeq TZ_{1}\oplus\operatorname{\mathcal{L}}|_{Z_{1}}$ which when combined with the multiplicativity of total Chern classes yields the adjunction formula
$c(TX|_{Z_{1}})=(1+c_{1}(\operatorname{\mathcal{L}}))c(Z_{1}).$
Expansion of the formula at $c_{2}(TX|_{Z_{1}})$ yields the second equality in the following computation:
$\chi(Z_{1})$ $=\int_{Z_{1}}c_{2}(Z_{1})=\int_{Z}c_{2}(TX|_{Z_{1}})-c_{1}(\operatorname{\mathcal{L}})c_{1}(X)+c_{1}(\operatorname{\mathcal{L}})^{2}$
$=\int_{X}c_{1}(\operatorname{\mathcal{L}})\left(c_{2}(X)-c_{1}(\operatorname{\mathcal{L}})c_{1}(X)+c_{1}(\operatorname{\mathcal{L}})^{2}\right).$
The first equality follows from the fact that $c_{2}(Z_{1})$ is the Euler class of $Z_{1}$. The third equality follows from $[Z_{1}]=\operatorname{PD}_{X}c_{1}(\operatorname{\mathcal{L}})$. Similar arguments provide us with
$c(TX|_{B})=(1+c_{1}(\operatorname{\mathcal{L}}))^{2}c(TX|_{B})\implies\chi(B)=\int_{X}c_{1}(\operatorname{\mathcal{L}})^{2}(c_{1}(X)-2c_{1}(\operatorname{\mathcal{L}})).$
Finally we apply $-\operatorname{Crit}\pi=\chi(X)-2\chi(Z_{1})+\chi(B)$ from Theorem 2.4.2 and add up the computed terms to obtain the general formula. In the Fano case with $\operatorname{\mathcal{L}}=A_{X}^{\otimes k}$, Equation (5.0.1) yields
$\int_{X}2c_{2}(X)c_{1}(\operatorname{\mathcal{L}})=\int_{X}2c_{2}(X)c_{1}(A_{X}^{\otimes k})=\int_{X}2kc_{2}(X)c_{1}(A_{X})=\int_{X}2kc_{2}(X)c_{1}(X)=48k.$
Substituting $c_{1}(\operatorname{\mathcal{L}})=kc_{1}(A_{X})$ into the remaining terms of the general formula then gives the second formula in the statement of the lemma. ∎
5.1. Classification of polarized $K3$ surfaces of low genus
Let $\operatorname{\mathcal{L}}\to X$ be very ample line bundle over a $\dim_{\mathbb{C}}X=3$ Fano variety with $\operatorname{\mathcal{L}}^{\otimes\operatorname{ind}_{X}}=A_{X}$. This implies that $\operatorname{\mathcal{L}}|_{Z}$ is very ample as well for $Z\subset X$ an anticanonical divisor. The following theorem reviews classification results of such polarized $K3$ surfaces $(Z,\operatorname{\mathcal{L}})$ by Saint-Donat [SD74] (for $g(Z,\operatorname{\mathcal{L}})\leq 5$) and Mukai [M87] (for $6\leq g(Z,\operatorname{\mathcal{L}})\leq 10$).
@@ 5.1. Classification of polarized $K3$ surfaces of low genus
$g=1+\frac{1}{2}\int_{Z}c_{1}(\operatorname{\mathcal{L}})^{2}=1+\frac{1}{2}\operatorname{ind}_{X}\int_{X}c_{1}(\operatorname{\mathcal{L}})^{3}=1+\frac{1}{2}\operatorname{ind}_{X}^{-2}\int_{X}c_{1}(A_{X})^{3}.$
Then $(Z,\operatorname{\mathcal{L}})$ has a projective model $(\mathcal{M}_{g}\subset\mathbb{CP}^{N_{g}},\vec{d})$ which is entirely determined by $g$.
For $g=3$, $Z$ is a $\deg=4$ hypersurface in $\mathbb{CP}^{3}$ with $\operatorname{\mathcal{L}}=\mathcal{O}_{1}$ so the model is $(\mathcal{M}_{3}=\mathbb{CP}^{4}\subset\mathbb{CP}^{4},\vec{d}=(1,4))$. For $g=4$, $Z$ is an intersection of $\deg=2$ and $\deg=3$ hypersurfaces in $\mathbb{CP}^{4}$ with $\operatorname{\mathcal{L}}=\mathcal{O}_{1}\rightarrow\mathbb{CP}^{4}$, so the model is $(\mathcal{M}_{5}=\mathbb{CP}^{5},\vec{d}=(1,2,3))$. For $g=5$, $Z$ is an intersection of three $\deg=2$ hypersurfaces in $\mathbb{CP}^{5}$ with $\operatorname{\mathcal{L}}=\mathcal{O}_{1}\rightarrow\mathbb{CP}^{5}$, so $(\mathcal{M}_{5}=\mathbb{CP}^{5},\vec{d}=(1,2,2,2))$. In the $6\leq g\leq 10$ cases, the $\mathcal{M}_{g}$ have positive codimension in $\mathbb{CP}^{N_{g}}$ and will be described in the following subsection.
@@ 5.2. Examples of pencil pairs from anticanonical divisors on Fano $3$-folds
Combining the above results with Theorems 4.1.3 and 4.2.2, we have the following theorem.
@@
Indeed, under the hypothesis of Theorem 5.2.1, Theorem 5.1.1 provides a projective model for an anticanonical divisor of each $X^{\pm}$ which are $K3$ surfaces. These projective models are entirely determined the genus, so the $(X^{\pm},A_{X^{\pm}})$ have divisors with a common projective model (Definition 4.1.2). By Lemma 4.1.3, there is a symplectomorphism $\phi=\phi_{1}$ for which the $(X^{\pm},A_{X^{\pm}},\phi)$ form a pencil pair. By Theorem 4.2.2, there are $\phi_{k}$ for which the $(X^{\pm},A_{X^{\pm}}^{\otimes k},\phi_{k})$ are pencil pairs for all $k\in\mathbb{Z}_{>0}$.
To find all examples of pencil pairs determinable from Theorem 5.2.1, we cross reference the above theorems with the complete classification of $\dim_{\mathbb{C}}=3$ Fano varieties by Iskovskih [I77, I77] and Mori-Mukai [MM82]. There are $105$ such families of varieties which are listed in the Fanography database [Be24] along with relevant data such as indices and Hodge diamonds (which of course determine Euler characteristics). [^1] See https://www.fanography.info/about for a complete list of resources from which the database is compiled.
We collect the relevant examples of our Fano $X$ into families (having at least two members) organized by their $\operatorname{ind}_{X}$ and $\int_{X}c_{1}(A_{X})^{3}$. Members of the families are labeled by their indices $a$-$b$ as in [Be24], with $a$ being the Picard rank. For each, a short description of $X$ is given. According to Theorem 5.2.1 any two members of a given family form a pencil pair.
5.2.1. Case $\operatorname{ind}_{X}=1,\int_{X}c_{1}(A_{X})^{3}=10\implies g(Z,A_{X})=6$
@@
    $X$ is an intersection of $\mathcal{M}_{6}=G(2,5)\subset\mathbb{CP}^{9}$ with two hyperplanes and a quadric for which $\chi(X)=-16$.
@@
    $X$ is a divisor of $\deg=(1,3)$ on $\mathbb{CP}^{1}\times\mathbb{CP}^{3}$ – that is, a blow up of $\mathbb{CP}^{3}$ along an intersection of two cubic hypersurfaces – having $\chi(X)=-14$.
Here is the projective model is $(\mathcal{M}_{6}\subset\mathbb{CP}^{9},\vec{d}=(1,1,1,2))$ following [M87]. We obtain Weinstein $4$-manifolds $(F_{k},\beta_{k})$ given by a generic $A_{X}^{\otimes k}$ divisor in either $X$, minus its intersection with another such generic divisor (that is the base locus of an associated pencil). The $F_{k}$ are smooth fibers of Lefschetz fibrations on $X$ minus a divisor for $A_{X}^{\otimes k}$ associated to the pencil for $A_{X}^{\otimes k}$. Since these divisors can be viewed as hyperplane sections in a projective variety, we can also view the $F_{k}$ as affine varieties.
Applying Theorem 5.2.1, we have factorizations of $\tau_{\partial F_{k}}$ into products of
$\chi+48k+10k^{2}(4k-3),\quad\chi=-16,-14$
Dehn twists along Lagrangian spheres in $F_{k}$.
5.2.2. Case $\operatorname{ind}_{X}=1,\int_{X}c_{1}(A_{X})^{3}=12\implies g(Z,A_{X})=7$
  1. (1-6)
    $X$ is an intersection of $\mathcal{M}_{7}\subset\mathbb{CP}^{15}$ of $\dim_{\mathbb{C}}\mathcal{M}=10$ with four hyperplanes, with $\chi(X)=-10$ as described in [M87].
  2. (2-5)
    $X$ is an intersection of $\deg=(0,3)$ and $\deg=(1,1)$ hypersurfaces in $\mathbb{CP}^{1}\times\mathbb{CP}^{4}$ (that is, a blow up of a cubic in $\mathbb{CP}^{4}$ along an intersection of two hyperplane sections) having $\chi(X)=-6$.
  3. (2-6)
    $X$ is a $\deg=(2,2)$ divisor in $\mathbb{CP}^{2}\times\mathbb{CP}^{2}$ having $\chi(X)=-12$.
  4. (3-1)
    $X$ is a double branched cover of $(\mathbb{CP}^{1})^{3}$ over a $\deg=(2,2,2)$ divisor having $\chi(X)=-8$.
In this case the projective model of [M87] is $(\mathcal{M}_{7}\subset\mathbb{CP}^{15},\vec{d}=(1,1,1,1,1))$.
$-\chi+48k+12k^{2}(4k-3),\quad\chi=-10,-6,-12,-8$
Dehn twists along Lagrangian spheres in $F_{k}$.
5.2.3. Case $\operatorname{ind}_{X}=1,\int_{X}c_{1}(A_{X})^{3}=14\implies g(Z,A_{X})=8$
  1. (1-7)
    $X$ is a the intersection of $\mathcal{M}_{8}=G(2,6)\subset\mathbb{CP}^{14}$ with nine hyperplanes with $\chi(X)=-6$.
  2. (2-7)
    $X$ is an intersection of $\deg=(0,2)$ and $\deg=(1,2)$ divisors on $\mathbb{CP}^{1}\times\mathbb{CP}^{4}$ with $\chi(X)=-4$.
  3. (2-8)
    Let $Y$ be the $\mathbb{CP}^{1}$-bundle given by projectivizing the $\mathbb{C}^{2}$-bundle $\mathbb{C}\oplus\mathcal{O}_{1}\rightarrow\mathbb{CP}^{2}$. Then $X$ is a double branched cover of $Y$ whose branch locus is an anticanonical divisor in $Y$ with $\chi(X)=-12$.
  4. (3-2)
    $X$ is a divisor in a $\mathbb{CP}^{2}$ bundle over $\mathbb{CP}^{1}\times\mathbb{CP}^{1}$ having $\chi(X)=2$.
$-\chi+48k+14k^{2}(4k-3),\quad\chi=-6,-4,-12,2$
@@ 5.2.3. Case $\operatorname{ind}_{X}=1,\int_{X}c_{1}(A_{X})^{3}=14\implies g(Z,A_{X})=8$
5.2.4. Case $\operatorname{ind}_{X}=1,\int_{X}c_{1}(A_{X})^{3}=16\implies g(Z,A_{X})=9$